# Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense

### v5

@@ Abstract
Deep learning technology has brought convenience and advanced developments but has become untrustworthy due to its sensitivity to adversarial attacks. Attackers may utilize this sensitivity to manipulate predictions. To defend against such attacks, existing anti-adversarial methods typically counteract adversarial perturbations post-attack, while we have devised a proactive strategy that preempts by safeguarding media upfront, effectively neutralizing potential adversarial effects before the third-party attacks occur. This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features. A forward-backward cascade learning algorithm is used to compute protective perturbations, starting with forward propagation optimization to achieve rapid convergence, followed by iterative backward propagation learning to alleviate overfitting. This strategy offers state-of-the-art transferability and protection across various systems. With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses. We have also devised the first, to our knowledge, effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised. This work provides a new direction to developing proactive defenses against adversarial attacks.
@@ 1 Introduction
Deep learning technologies have been demonstrated to be sensitive to adversarial attacks [18]. An adversarial attacker can utilize this weakness to manipulate medias such as images, audios, and videos. For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access the face authentication system [47]. The attacker can also add slight noise on traffic signs, possibly causing an accident during autonomous driving [15]. Defending against such attacks is a critical but unsolved challenge. Passive defenses [6, 29, 34, 22], on the one hand, trade off classification accuracy for resistance against threats [54], but may be vulnerable to counterattacks [7, 46]. Proactive defenses [32, 17], on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications. Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is untransferable to other systems. Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [39].
@@
To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Fig. 1. Fast Preemption preemptively “attacks” (manipulates for enhancement) the medias from the user side before being attacked by the third party. Specifically, Fast Preemption learns robust features and predicts possible adversarial perturbations, then enhances these robust features and reverses the predicted perturbations for purpose of protection. In this case, when these protected medias are further attacked, these adversarial perturbations will be automatically neutralized. This strategy has three distinct advantages. (i) A proactive preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation for the classification system. (ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system [16, 39]. (iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated within a classification model) is breached, all inputs are affected. In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings. Therefore, Fast Preemption meets four basic requirements:
@@
We propose the Fast Preemption defense includes three components: a classifier model for automatically labeling inputs, a backbone model for learning features (differ from the classifier), and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations. We conducted evaluation on two datasets [26, 2] across six models [19, 6, 35, 29, 32, 13], defending against two typical adversarial attacks [31, 11]. The results demonstrate the superior performance of the Fast Preemption strategy in terms of clean and robust accuracies, transferability, and time cost compared with those of benchmark training-time, test-time, and preemptive adversarial defenses [6, 29, 34, 32, 17]. For example, in an end-to-end comparison with the state-of-the-art (SOTA) preemptive defense [32] on the ImageNet dataset [2], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [11] and was 33 times faster.
@@
Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [43].” On the basis of that assumption, Bi-Level [29] first proposed a reversion attack but failed, claiming their defense was irreversible. However, our reversion attack successfully defeated their defense, establishing ours as the first effective adaptive reversion attack in the white-box setting, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible. Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense. Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier). If clean accuracy increases after reversion, the reversion succeeded. If it decreases after reversion (due to distortion), reversion failed.
@@
The results from reversion attack testing demonstrate that the Preemptive Reversion attack can smartly erase protective perturbations when all defense details are compromised (i.e., a white-box situation), whereas blind JPEG distortion and adversarial purification [34] only further distort an image. However, this white-box situation is impractical with the Fast Preemption strategy for real-world applications because the proposed framework enables users to individually customize their defense settings. Apart from a white-box situation, Fast Preemption protection cannot be reversed using existing reversion approaches.
@@
We summarize our contributions as follows: (i) We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module. It is the first to use different models for the classifier and backbone. (ii) We present a forward-backward (Fore-Back) cascade learning algorithm that speeds up preemptive defense for efficient user interaction. (iii) We present the first effective adaptive reversion attack (Preemptive Reversion) for evaluating the reliability of preemptive defenses in a white-box setting and a more comprehensive protocol to evaluate whether reversion succeeds.
2 Related Work
@@
Existing adversarial defenses can be categorized into three primary types: training-time, test-time, and preemptive defenses. Appendix A provides a more detailed review.
@@
Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels [18, 42, 41, 19, 13, 50, 14, 35, 54, 6, 29, 44]. However, they have two shortcomings. First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness. Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy.
@@
Test-time defenses are divided to detection and purification. Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs [9, 56, 27, 22]. However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks. Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy. In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations. Earlier studies used image distortion for purification [48, 20, 37]. Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious [38, 34]. The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [7, 46].
Preemptive defenses seek to generate robust examples (REs) that replace the original images, which are subsequently discarded [40, 32, 17]. These REs are inherently more difficult to manipulate, thereby providing resistance against adversarial attacks. However, such approaches face significant limitations, including high computational costs, limited transferability across models, and, in some cases, lower robustness compared to other defense mechanisms.
@@ 3.3 Forward-Backward Cascade Learning
The Fore-Back cascade learning algorithm is used to compute protective perturbations. The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth. We define this as forward propagation ($\rightarrow$). Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models. Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation. Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability. In summary, the forward and backward propagation both enhance robustness, but in different optimization directions. The complete algorithm is presented in Algorithm 1 of Appendix B.
@@ 3.3.2 Preliminary - Backward Propagation
where $\epsilon$ is the setting of the maximum distance and $\alpha$ is the step size. The final output of Eq. 8 is AE $x_{t}^{a}$. Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next robust example, $x_{t+1}$:
@@
The SOTA preemptive defense strategy [32] conducts only backward propagation learning. However, adversarial attacks can be randomly initialized, so backward propagation requires many steps to converge (minimum ten steps). The Fast Preemption strategy accelerates defense by running forward propagation before running backward propagation learning.
@@ 3.3.3 Combining Forward and Backward Propagation
the final output of which is robust example $x_{t}^{r}$. Subsequently, protective perturbations $x_{t}^{r}-x_{t}$ are directly added to $x_{t}$ in accordance with learning rate $\eta$:
@@ 4 Adaptive Attack: Preemptive Reversion
A reversion attack is used to evaluate whether a preemptive defense is reversible. While Bi-Level [32] first proposed a reversion attack but failed, claiming their defense was irreversible, our reversion attack successfully defeated their defense, establishing ours as the first effective reversion attack. In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks. Preemptive Reversion is the only feasible approach to reverse preemptive defense in the white-box scenario. Nevertheless, without full knowledge of the backbone model and settings, our Fast Preemption defense is hard to reverse, so it is reliable.
@@ 4.1 Algorithm
When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Fig. 5). On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig. 3. The complete algorithm is presented in Algorithm 2 of Appendix B. Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$. Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
@@ 4.2 Attack Scenarios
We define two scenarios for a reversion attack in accordance with the knowledge the attacker has about the backbone model and defense algorithm:
  • White-Box: The attacker knows both the backbone model and defense algorithm (including settings), so the attacker can conduct a secondary defense the same as the user does. This is the strongest, yet impractical, assumption because different users may use different settings.
  • Black-Box: The attacker knows the defense algorithm (typically public) but not the backbone model (infinite selections for users). The challenge for the attacker in this scenario is that the crucial features learned in the secondary defense may differ from those learned by the user.
Our proposed Preemptive Reversion attack is the first effective reversion attack for evaluating the reversibility of preemptive defenses, including our Fast Preemption defense, but it is feasible only in the white-box scenario. Without full knowledge of the defense, Fast Preemption is reliable.
@@ 4.3 Evaluation Protocol
If a reversion attack succeeds, the improved clean accuracy should drop. However, image distortion can create a similar effect. Consequently, a drop in clean accuracy does not necessarily indicate the success of a reversion attack. To address this issue, we proposed a comprehensive protocol that can distinguish successful reversion from distortion. First, we purposely give incorrect labels to partial inputs to guarantee that the clean accuracy after being defended is lower than the original accuracy. In this situation, if the clean accuracy approaches close to the original accuracy, the reversion attack succeeded. Otherwise, the reversion attack failed, and the deterioration was caused by distortion.
5 Experiments
5.1 Datasets, Classifier, and Backbone Models
We conducted experiments on two datasets: CIFAR-10 [26] and ImageNet [2]. For CIFAR-10, ResNeXt29-32x4d [21] and PreActResNet-18 [19] were selected as the classifier and backbone model, respectively. For ImageNet, DeiT Base [45] and ConvNeXt-L+ConvStem [44] were selected as the classifier and backbone model, respectively. The model architectures are listed in Tab. 4 of Appendix D.
5.2 Attacks and Defenses
Adversarial attacks. We evaluated defenses against two adversarial attacks: PGD [31] and AutoAttack [11]. For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, and FAB [10] and Square Attack [4]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [5]). Reversion attacks. We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective. Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [34]).
Training-time defenses: We compared Fast Preemption with WideResNet-94-16 [6] and Swin-L [29] on the CIFAR-10 and ImageNet datasets, respectively. The model architectures of these defenses differ from the backbone models. Test-time defenses: We used DiffPure [34] as the benchmark test-time defense. The inputs were purified by DiffPure and then fed into WideResNet-94-16 [6] or Swin-L [29] in accordance with the dataset used. Preemptive defenses. We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [32], and the latest preemptive defense, $A^{5}$ [17].

### v6

@@ Abstract
Deep learning has advanced significantly but remains vulnerable to adversarial attacks, compromising its reliability. While conventional defenses typically mitigate perturbations post-attack, few studies explore proactive strategies for preemptive protection. This paper proposes Fast Preemption, a novel defense that neutralizes adversarial effects before third-party attacks. By employing distinct models for input labeling and feature extraction, Fast Preemption enables an efficient, transferable preemptive defense with state-of-the-art robustness across diverse systems. To further enhance efficiency, we introduce a forward-backward cascade learning algorithm that generates protective perturbations, leveraging forward propagation for rapid convergence and iterative backward propagation to mitigate overfitting. Executing in just three iterations, Fast Preemption surpasses existing training-time, test-time, and preemptive defenses. Additionally, we propose the first effective white-box adaptive reversion attack to evaluate the reversibility of preemptive defenses, demonstrating that our approach remains secure unless the backbone model, algorithm, and settings are entirely compromised. This work establishes a new paradigm for proactive adversarial defense.
@@ 1 Introduction
Deep learning models are highly susceptible to adversarial attacks [18], enabling adversaries to manipulate media such as images, audio, and video. For instance, an attacker can alter a facial image’s identity without perceptible changes, bypassing facial authentication systems [47], or introduce imperceptible noise to traffic signs, potentially compromising autonomous driving safety [15]. Defending against such attacks remains a critical yet unresolved challenge. Passive defenses [6, 29, 34, 22] enhance robustness at the cost of classification accuracy [54] and remain susceptible to adaptive attacks [7, 46]. Proactive defenses [32, 17], while mitigating threats preemptively, often lack real-time efficiency and require paired protection-inference modules, limiting transferability. Moreover, users relying on third-party models must contend with potential backdoor threats [39].
@@
To address these challenges, we propose a proactive defense strategy, Fast Preemption, as illustrated in Fig. 1. Fast Preemption preemptively “attacks” media on the user side to enhance robustness before potential third-party attacks. Specifically, it learns robust features, predicts possible adversarial perturbations, and then amplifies these features while reversing the predicted perturbations for protection. As a result, any subsequent adversarial attack is inherently neutralized. This approach offers three key advantages. (i) Unlike passive defenses, Fast Preemption enhances threat resistance without compromising classification accuracy and requires no modifications to the classification system. (ii) It supports plug-and-play deployment without paired protection-inference modules, mitigating security and privacy concerns associated with external model providers [16, 39]. (iii) Traditional training- and test-time defenses, integrated within classification models, risk full-system compromise if breached. In contrast, Fast Preemption localizes vulnerabilities to individual users, as each employs distinct configurations. Thus, Fast Preemption satisfies four fundamental requirements:
@@
We introduce Fast Preemption, a preemptive defense comprising three key components: a classifier model for automatic input labeling, a backbone model (distinct from the classifier) for feature learning, and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations. We evaluate Fast Preemption on two datasets and thirteen models, defending against five representative adversarial attacks. Results demonstrate its superiority in clean and robust accuracy, transferability, and computational efficiency over benchmark training-time, test-time, and preemptive defenses [6, 29, 34, 32, 17]. For example, on CIFAR-10 [26], Fast Preemption outperformed the state-of-the-art (SOTA) preemptive defense [32], achieving 2.5% and 5.1% higher clean and robust accuracy, respectively, against AutoAttack (Standard) [11], while being 700 times faster.
@@
Shannon famously stated, “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them” [43]. Following this principle, Bi-Level [32] first introduced a reversion attack but failed, asserting their defense was irreversible. In contrast, our reversion attack successfully circumvented their defense, making ours the first effective adaptive reversion attack in the white-box setting, termed Preemptive Reversion, to assess the reversibility of preemptive defenses. Preemptive Reversion applies a secondary preemptive defense to the already protected example and then reverses the newly introduced perturbations. Since the secondary defense generates perturbations similar to the first under identical conditions, this process effectively tests a defense’s resilience to reversion. Additionally, we propose a more rigorous evaluation protocol, which mandates a decrease in clean accuracy after preemptive defense (achieved by selecting a suboptimal classifier). If clean accuracy improves post-reversion, the attack is successful; if it further deteriorates due to distortion, reversion has failed.
@@
Results of reversion attacks show that Preemptive Reversion effectively removes protective perturbations in a white-box setting, where all defense details are exposed. In contrast, blind JPEG distortion and adversarial purification [34] further degrade the image rather than neutralizing protections. However, this white-box assumption is unrealistic for real-world applications of Fast Preemption, as the framework allows users to customize their defense settings individually. Beyond the white-box scenario, existing reversion methods fail to circumvent Fast Preemption’s protection.
@@
We summarize our contributions as follows: (i) We introduce Fast Preemption, a novel preemptive defense that enhances transferability without requiring ground truths or paired modules. It is the first approach to utilize distinct models for the classifier and backbone. (ii) We propose the forward-backward (Fore-Back) cascade learning algorithm, which accelerates preemptive defense, enabling efficient user interaction. (iii) We develop Preemptive Reversion, the first effective adaptive reversion attack for assessing the robustness of preemptive defenses in a white-box setting. Additionally, we establish a more comprehensive evaluation protocol to determine reversion success.
2 Related Works
@@
Existing adversarial defenses can be classified into three main categories: training-time, test-time, and preemptive defenses. A detailed review is provided in Appendix F.
@@
Training-time defenses aim to develop robust deep learning models that can correctly classify adversarial examples (AEs) based on their ground truth labels [18, 42, 41, 19, 13, 50, 14, 35, 54, 6, 29, 44]. However, they have two key limitations. First, they require large amounts of high-quality, diverse data for training. Second, training-time defenses often sacrifice clean accuracy to improve robust accuracy.
@@
Test-time defenses are typically divided into detection and purification. Detection models identify AEs by distinguishing them from benign inputs [9, 56, 27, 22], but detection is challenging for unseen attacks, as detectors are often trained on known AEs. Furthermore, detection may misclassify benign images as AEs, reducing clean accuracy. In contrast, adversarial purification aims to remove perturbations in AEs. Early approaches used image distortion for purification [48, 20, 37], while recent methods generate benign replacements regardless of input type [38, 34]. A key limitation of purification is that the difference between original and generated examples can degrade clean accuracy, and counterattack potential remains [7, 46].
@@
Preemptive defenses aim to generate robust examples (REs) that replace the original images, which are then discarded [40, 32, 17]. These REs are inherently more resilient to manipulation, offering resistance against adversarial attacks. However, these approaches face significant limitations, including high computational costs, limited transferability across models, and, in some cases, reduced robustness compared to other defense mechanisms.
Alfarra et al. [3] proposed a test-time defense incorporating preemptive protection as an anti-adversary layer guided by a standard classifier. However, their method is inherently vulnerable if the classifier is compromised, a limitation they explicitly acknowledge. This weakness undermines its practical viability in test-time scenarios. While they describe this setting as the “least realistic,” we argue that it aligns with the standard white-box attack model, where adversaries have full access to the classifier. Under even stricter assumptions, adversaries may also be aware of the defense mechanism itself (white-box reversion). Moreover, their approach relies solely on forward propagation learning, increasing susceptibility to overfitting.
@@ 3.3 Forward-Backward Cascade Learning
The Fore-Back cascade learning algorithm is used to compute protective perturbations. The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth. We define this as forward propagation ($\rightarrow$). Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models. Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation. Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability. In summary, the forward and backward propagation both enhance robustness, but in different optimization directions. The complete algorithm is presented in Algorithm 1 of Appendix D.
@@ 3.3.2 Preliminary - Backward Propagation
where $\epsilon$ is the setting of the maximum distance and $\alpha$ is the step size. The final output of Eq. 8 is AE $x_{t}^{a}$. Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next RE, $x_{t+1}$:
@@
The SOTA preemptive defense strategy [32] relies solely on backward propagation learning, which requires multiple steps to converge due to the random initialization of adversarial attacks (at least ten steps). In contrast, the Fast Preemption strategy accelerates defense by incorporating forward propagation before backward propagation learning.
@@ 3.3.3 Combining Forward and Backward Propagation
the final output of which is RE $x_{t}^{r}$. Subsequently, protective perturbations $x_{t}^{r}-x_{t}$ are directly added to $x_{t}$ in accordance with learning rate $\eta$:
@@ 4 Adaptive Attack: Preemptive Reversion
A reversion attack is used to assess the reversibility of a preemptive defense. While Bi-Level [32] first proposed a reversion attack but failed to reverse their defense, claiming it was irreversible, our reversion attack successfully overcame their defense, making ours the first effective reversion attack. In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks. Preemptive Reversion is the only viable approach to reverse preemptive defenses in the white-box scenario. However, without full knowledge of the backbone model and settings, our Fast Preemption defense remains difficult to reverse, ensuring its reliability.
@@ 4.1 Algorithm
When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Fig. 5). On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig. 3. The complete algorithm is presented in Algorithm 2 of Appendix D. Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$. Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
@@ 4.2 Attack Scenarios
We define two reversion attack scenarios via the attacker’s knowledge of the backbone model and defense algorithm:
  • White-Box: The attacker has complete knowledge of both the backbone model and defense algorithm (including settings), enabling them to replicate the secondary defense. This scenario, while the strongest, is impractical due to varying user settings.
  • Black-Box: The attacker knows the defense algorithm (typically public) but lacks knowledge of the backbone model, which can vary across users. The challenge in this scenario is that the features learned in the secondary defense may differ from those learned by the user.
Our proposed Preemptive Reversion attack is the first effective method for evaluating the reversibility of preemptive defenses, including Fast Preemption, but it is only feasible in the white-box scenario. Without full knowledge of the defense, Fast Preemption remains reliable.
@@ 4.3 Evaluation Protocol
If a reversion attack succeeds, clean accuracy should decrease. However, image distortion can lead to a similar outcome, meaning a drop in clean accuracy does not necessarily indicate a successful reversion. To address this, we propose a comprehensive protocol to distinguish successful reversion from distortion. First, we intentionally assign incorrect labels to partial inputs, ensuring the clean accuracy after defense is lower than the original. If clean accuracy approaches the original value, the reversion attack is successful; otherwise, the drop in accuracy is due to distortion.
5 Experiment Settings
Datasets, Classifier, and Backbone Models. We conducted experiments on two datasets: CIFAR-10 [26] and ImageNet [2]. For CIFAR-10, ResNeXt29-32x4d [21] and PreActResNet-18 [19] were selected as the classifier and backbone model, respectively. For ImageNet, DeiT Base [45] and ConvNeXt-L+ConvStem [44] were selected as the classifier and backbone model, respectively. The model architectures are listed in Tab. 9 of Appendix E.
Adversarial attacks. We evaluated defenses against five adversarial attacks: white-box PGD [31], AutoAttack [11], and Carlini-Wagner (CW) attack [7]; black-box Deformation-Constrained Warping (DeCoW) [28] and Meta Conditional Generator (MCG) [53]. For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, FAB [10], and Square Attack [4]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [5]).
Reversion attacks. We examined: (i) whether protected samples induced by preemptive defenses cannot be easily reversed to their original unprotected state, which we define as “irreversible”; and (ii) the effectiveness of the proposed adaptive reversion attack, Preemptive Reversion. Along with Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [34]).
Training-time defenses. We compared Fast Preemption with WideResNet-94-16 [6] and Swin-L [29] on the CIFAR-10 and ImageNet datasets, respectively. The architectures of these models differ from the backbone models. To assess the transferability of preemptive defenses, we conducted experiments across additional nine victim models, with results presented in Tab. 7 of Appendix C.
Test-time defenses. We used DiffPure [34] as the benchmark test-time defense. The inputs were purified by DiffPure and then fed into WideResNet-94-16 [6] or Swin-L [29] depending on the dataset.
Preemptive defenses. We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [32], and the latest preemptive defense, $A^{5}$ [17].